# ORCID: $0$-0002-6512-4229 (Om Prakash)

### v1

@@ 3 $(\theta,\Im)$-cyclic codes over $\mathbb{F}_{q}$
In the following, we discuss the main algebraic properties of $(\theta,\Im)$-cyclic codes in $R=\mathbb{F}_{q}[x;\theta,\Im]$ and provide a necessary and sufficient condition for these codes to contain their Euclidean duals. In MH , Boulagouaz and Leroy introduced the notion of $(f,\gamma,\Delta)$-cyclic codes. Moreover, a $(\theta,\Im)$-cyclic code $\mathcal{C}$ is the subset of $\mathbb{F}_{q}^{n}$ consisting of the coordinates of the elements of $Rg(x)/\langle x^{n}-1\rangle$ in the basis $\{1,x,\dots,x^{n-1}\}$ for some right monic factors $g(x)$ of $x^{n}-1$.
Theorem 3.1
@@
    A $(\theta,\Im)$-cyclic code of length $n$ corresponding to $Rg(x)/\langle x^{n}-1\rangle$ is a free left $\mathbb{F}_{q}$-module of dimension $n-r$ where $r=\deg~{}g(x)$.
@@
    $T_{\theta,\Im}^{k}(g_{0},g_{1},\dots,g_{r},0,0,\dots,0),~{}~{}~{}~{}\text{for}~{}0\leq k\leq n-r-1.$
Proof
@@
    We have $x^{n}-1=h(x)g(x)$ for some monic polynomials $h(x)\in R$. Hence, as left $R$-module, we have $Rg(x)/\langle x^{n}-1\rangle\cong R/\langle h(x)\rangle$. Since $h$ is monic, $R/\langle h(x)\rangle$ is a free left $\mathbb{F}_{q}$-module of rank $deg~{}h(x)=n-r$.
@@
    We have $T_{\theta,\Im}^{k}(v_{0},v_{1},\dots,v_{n-1})\in\mathcal{C}$ for any $k\geq 0$. On the other hand, it is clear that $g,xg,x^{2}g,\dots,\\ x^{n-r-1}g$ are left linearly independent over $\mathbb{F}_{q}$, all are taken modulo $x^{n}-1$ and hence form a basis of $Rg(x)/\langle x^{n}-1\rangle$. In codewords representation, this implies that the vectors $T_{\theta,\Im}^{k}(g_{0},g_{1},\dots,g_{r},0,\dots,0)$ form a left $\mathbb{F}_{q}$-basis for $\mathcal{C}$, $10\leq k\leq n-r-$ 1 .
Theorem 3.2
@@
Proof
Let $g(x)\in\mathcal{C}$ be a monic smallest degree polynomial among nonzero polynomials in $\mathcal{C}$ and $c(x)\in\mathcal{C}$. Then by Theorem 2.1, there exist unique polynomials $q(x)$ and $r(x)$ in $R$ such that $c(x)=q(x)g(x)+r(x)$ where $r(x)=0~{}\text{or}~{}\deg r(x)<\deg g(x).$ As $\mathcal{C}$ is a left $R$-submodule, we have $r(x)=c(x)-q(x)g(x)\in\mathcal{C}$. This is a contradiction to the assumption that $g(x)$ is of the smallest degree in $\mathcal{C}$ unless $r(x)=0$. This implies $c(x)=q(x)g(x)$ and hence $\mathcal{C}$ is a $(\theta,\Im)$-cyclic submodule generated by $g(x)$.
Theorem 3.3
@@
Proof
Consider a monic smallest degree polynomial $g(x)$ in $\mathcal{C}$. From Theorem 2.1, there exist polynomials $q(x)$ and $r(x)$ in $R$ such that $x^{n}-1=q(x)g(x)+r(x)$, where $\deg r(x)<\deg g(x).$ Since $g(x)$ and $x^{n}-1=0$ are in $\mathcal{C}$, this implies $r(x)=(x^{n}-1)-q(x)g(x)\in\mathcal{C}$. But, $g(x)$ is smallest in $\mathcal{C}$. Therefore, $r(x)=0$ and hence $g(x)$ is a right divisor of $x^{n}-1$.
Let $\mathcal{C}$ be a $(\theta,\Im)$-cyclic code of length $n$ over $\mathbb{F}_{q}$ generated by the right divisor $g(x)$ of $x^{n}-1$, where $g(x)=g_{0}+g_{1}x+\cdots+g_{r}x^{r}\in R$ and $g_{r}=1$. Then from the above discussion, we can conclude that $\mathcal{C}$ is a free left $\mathbb{F}_{q}$-module of dimension $k=n-\deg g(x)$. Now, by using (Leroy, , Theorem 3.2), the generator matrix of $\mathcal{C}$ is given by
@@
$\tau:~{}~{}~{}~{}\mathbb{F}_{q}^{n}~{}~{}~{}~{}~{}\longrightarrow~{}~{}~{}~{}\mathbb{F}_{q}[x;\theta,\Im]/\langle x^{n}-1\rangle$
@@
$c(x)$ $=a(x)g(x)~{}\text{for some}~{}a(x)\in R$
@@
$=a(x)(x^{n}-1)=0~{}\text{in}~{}\mathbb{F}_{q}[x;\theta,\Im]/\langle x^{n}-1\rangle.$
@@
This implies that $c(x)=q(x)g(x)\in\langle g(x)\rangle=\mathcal{C}$ as $h^{\prime}(x)$ is a nonzero polynomial.
@@
$\tau(c(T_{\theta,\Im})(h^{\prime}))=c(x)h^{\prime}(x)=a(x)g(x)h^{\prime}(x)=a(x)h(x)g(x)=a(x)(x^{n}-1)=0.$
for some $a(x)$ in $\mathbb{F}_{q}[x;\theta,\Im]/\langle x^{n}-1\rangle$ and $c(x)h^{\prime}(x)$ is taken modulo $x^{n}-1$. This implies $c(T_{\theta,\Im})(h^{\prime})=0$. Thus, $0=c(T_{\theta,\Im})(h^{\prime})=c_{0}+c_{1}T_{\theta,\Im}(h^{\prime})+c_{2}T_{\theta,\Im}^{2}(h^{\prime})+\cdots+c_{n-1}T_{\theta,\Im}^{n-1}(h^{\prime}).$ This shows that $(c_{0},c_{1},c_{2},\dots,c_{n-1}).H=0$ for any $c=(c_{0},c_{1},c_{2},\dots,c_{n-1})\in\mathcal{C}$. Also, $\tau(T_{\theta,\Im}^{k}(h^{\prime}))=x^{k}h^{\prime}(x)$ for $k=0,1,\dots,n-deg~{}h^{\prime}(x)-1=deg~{}g(x)-1$ and hence $\{h^{\prime},T_{\theta,\Im}(h^{\prime}),T_{\theta,\Im}^{2}(h^{\prime}),\dots,T_{\theta,\Im}^{r-1}(h^{\prime})\}$ are linearly independent.
@@
As $h^{\prime}(x)$ is a nonzero polynomial, we have $a(x)-c(x)b(x)g(x)=0$, which gives $a(x)=c(x)b(x)g(x)$. Therefore, $a(x)\in\mathcal{C}=\langle g(x)\rangle$. Thus, $\mathcal{C}^{\perp}\subseteq\mathcal{C}$.
@@ 4 $(\gamma,\Delta)$-cyclic codes over $\mathscr{R}_{q,s}$
In this section, our main focus is on the algebraic properties of $(\gamma,\Delta)$-cyclic codes over $\mathscr{R}_{q,s}$ via decomposition over $\mathbb{F}_{q}$. To do so, we consider a linear code $\mathcal{C}$ of length $n$ over $\mathscr{R}_{q,s}$. Now, we define
$\mathcal{C}_{i}=\left\{t_{i}\in\mathbb{F}_{q}^{n}~{}|~{}\sum_{i=0}^{s}\zeta_{i}t_{i}\in\mathcal{C},~{}\text{for some}~{}t_{0},t_{1},\dots,t_{i-1},t_{i+1},\dots,t_{s}\in\mathbb{F}_{q}^{n}\right\}$
for $0\leq i\leq s$. Then $\mathcal{C}_{i}$ is linear code of length $n$ over $\mathbb{F}_{q}$ and $\mathcal{C}$ can be decomposed as follows
@@
Hence, $\Delta$ is a $\gamma$-derivation of $\mathscr{R}_{q,s}$.
@@
Therefore, $\mathcal{C}$ is a $(\gamma,\Delta)$-cyclic code of length $n$ over $\mathscr{R}_{q,s}$.
Theorem 4.3
Let $\mathcal{C}=\displaystyle\bigoplus_{i=0}^{s}\zeta_{i}\mathcal{C}_{i}$ be a $(\gamma,\Delta)$-cyclic code of length $n$ over $\mathscr{R}_{q,s}$. Then $\mathcal{C}=\langle\zeta_{0}g_{0}(x),\\ \zeta_{1}g_{1}(x),\dots,\zeta_{s}g_{s}(x)\rangle$ and $|\mathcal{C}|=q^{(s+1)n-\sum_{i=0}^{s}\deg(g_{i}(x))}$, where $g_{i}(x)$ is a generator polynomial of $\mathcal{C}_{i}$ for $i=0,1,2,\dots,s$.
Proof
@@
On the other hand, we consider $\zeta_{0}f_{0}(x)g_{0}(x)+\zeta_{1}f_{1}(x)g_{1}(x)+\cdots+\zeta_{s}f_{s}(x)g_{s}(x)\in\langle\zeta_{0}g_{0}(x),\\ \zeta_{1}g_{1}(x),\dots,\zeta_{s}g_{s}(x)\rangle\subseteq\mathscr{R}_{q,s}[x;\gamma,\Delta]/\langle x^{n}-1\rangle$ where $f_{i}(x)\in\mathscr{R}_{q,s}[x;\gamma,\Delta]/\langle x^{n}-1\rangle$ for $i=0,1,2,\dots,s$. Then there exists $s_{i}(x)\in\mathbb{F}_{q}[x;\theta,\Im]/\langle x^{n}-1\rangle$ such that $\zeta_{i}f_{i}(x)=\zeta_{i}s_{i}(x)$ for $i=0,1,2,\dots,s$. This implies that $\langle\zeta_{0}g_{0}(x),\zeta_{1}g_{1}(x),\dots,\zeta_{s}g_{s}(x)\rangle\subseteq\mathcal{C}$. Thus, $\mathcal{C}=\langle\zeta_{0}g_{0}(x),\zeta_{1}g_{1}(x),\dots,\zeta_{s}g_{s}(x)\rangle$.
Moreover, $|\mathcal{C}|=|\mathcal{C}_{0}||\mathcal{C}_{1}|\cdots|\mathcal{C}_{s}|=q^{n-\deg(g_{0}(x))}q^{n-\deg(g_{1}(x))}\cdots q^{n-\deg(g_{s}(x))}\\ =q^{(s+1)n-\sum_{i=0}^{s}\deg(g_{i}(x))}$.
Theorem 4.4
@@
Conversely, let $\mathcal{C}^{\perp}\subseteq\mathcal{C}$ , then $\bigoplus_{i=0}^{s}\zeta_{i}\mathcal{C}_{i}^{\perp}\subseteq\displaystyle\bigoplus_{i=0}^{s}\zeta_{i}\mathcal{C}_{i}$. Now, considering modulo $\zeta_{i}$, we get $\mathcal{C}_{i}^{\perp}\subseteq\mathcal{C}_{i}$ for $i=0,1,2,\dots,s$. Thus, $h_{i}^{\prime}(x)h_{i}^{\prime}(x)$ is divisible by $x^{n}-1$ on the right for $i=0,1,2,\dots,s$.

### v2

@@ 3 $(\theta,\Im)$-cyclic codes over $\mathbb{F}_{q}$
This section presents the algebraic properties of $(\theta,\Im)$-cyclic codes in $R=\mathbb{F}_{q}[x;\theta,\Im]$ and provide a necessary and sufficient condition for these codes to contain their Euclidean duals. In [9], Boulagouaz and Leroy introduced the notion of $(f,\gamma,\Delta)$-cyclic codes. Moreover, a $(\theta,\Im)$-cyclic code $\mathcal{C}$ is the subset of $\mathbb{F}_{q}^{n}$ consisting of the coordinates of the elements of $Rg(x)/\langle x^{n}-1\rangle$ in the basis $\{1,x,\dots,x^{n-1}\}$ for some right monic factors $g(x)$ of $x^{n}-1$.
Theorem 3.1.
@@
    A $(\theta,\Im)$-cyclic code of length $n$ corresponding to $Rg(x)/\langle x^{n}-1\rangle$ is a free left $\mathbb{F}_{q}$-module of dimension $n-r$ where $r=\deg\leavevmode\nobreak\ g(x)$.
@@
    $T_{\theta,\Im}^{k}(g_{0},g_{1},\dots,g_{r},0,0,\dots,0),\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \text{for}\leavevmode\nobreak\ 0\leq k\leq n-r-1.$
Proof.
@@
    We have $x^{n}-1=h(x)g(x)$ for some monic polynomials $h(x)\in R$. Hence, as left $R$-module, we have $Rg(x)/\langle x^{n}-1\rangle\cong R/\langle h(x)\rangle$. Since $h$ is monic, $R/\langle h(x)\rangle$ is a free left $\mathbb{F}_{q}$-module of rank $deg\leavevmode\nobreak\ h(x)=n-r$.
@@
    We have $T_{\theta,\Im}^{k}(v_{0},v_{1},\dots,v_{n-1})\in\mathcal{C}$ for any $k\geq 0$. On the other hand, it is clear that $g,xg,x^{2}g,\dots,x^{n-r-1}g$ are left linearly independent over $\mathbb{F}_{q}$, all are taken modulo $x^{n}-1$ and hence form a basis of $Rg(x)/\langle x^{n}-1\rangle$. In codewords representation, this implies that the vectors $T_{\theta,\Im}^{k}(g_{0},g_{1},\dots,g_{r},0,\dots,0)$ form a left $\mathbb{F}_{q}$-basis for $\mathcal{C}$, $10\leq k\leq n-r-$ 1 .
Theorem 3.2.
@@
Proof.
Let $g(x)\in\mathcal{C}$ be a monic smallest degree polynomial among nonzero polynomials in $\mathcal{C}$ and $c(x)\in\mathcal{C}$. Then by Theorem 2.1, there exist unique polynomials $q(x)$ and $r(x)$ in $R$ such that $c(x)=q(x)g(x)+r(x)$ where $r(x)=0\leavevmode\nobreak\ \text{or}\leavevmode\nobreak\ \deg r(x)<\deg g(x).$ As $\mathcal{C}$ is a left $R$-submodule, we have $r(x)=c(x)-q(x)g(x)\in\mathcal{C}$. This is a contradiction to the assumption that $g(x)$ is of the smallest degree in $\mathcal{C}$ unless $r(x)=0$. This implies $c(x)=q(x)g(x)$ and hence $\mathcal{C}$ is a $(\theta,\Im)$-cyclic submodule generated by $g(x)$. ∎
Theorem 3.3.
@@
Proof.
Consider a monic smallest degree polynomial $g(x)$ in $\mathcal{C}$. From Theorem 2.1, there exist polynomials $q(x)$ and $r(x)$ in $R$ such that $x^{n}-1=q(x)g(x)+r(x)$, where $\deg r(x)<\deg g(x).$ Since $g(x)$ and $x^{n}-1=0$ are in $\mathcal{C}$, this implies $r(x)=(x^{n}-1)-q(x)g(x)\in\mathcal{C}$. But, $g(x)$ is smallest in $\mathcal{C}$. Therefore, $r(x)=0$ and hence $g(x)$ is a right divisor of $x^{n}-1$. ∎
Let $\mathcal{C}$ be a $(\theta,\Im)$-cyclic code of length $n$ over $\mathbb{F}_{q}$ generated by the right divisor $g(x)$ of $x^{n}-1$, where $g(x)=g_{0}+g_{1}x+\cdots+g_{r}x^{r}\in R$ and $g_{r}=1$. Then from the above discussion, we can conclude that $\mathcal{C}$ is a free left $\mathbb{F}_{q}$-module of dimension $k=n-\deg g(x)$. Now, by using [29, Theorem 3.2], the generator matrix of $\mathcal{C}$ is given by
@@
$\tau:\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathbb{F}_{q}^{n}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \longrightarrow\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathbb{F}_{q}[x;\theta,\Im]/\langle x^{n}-1\rangle$
@@
$c(x)$ $=a(x)g(x)\leavevmode\nobreak\ \text{for some}\leavevmode\nobreak\ a(x)\in R$
@@
$=a(x)(x^{n}-1)=0\leavevmode\nobreak\ \text{in}\leavevmode\nobreak\ \mathbb{F}_{q}[x;\theta,\Im]/\langle x^{n}-1\rangle.$
@@
This implies that $c(x)=q(x)g(x)\in\langle g(x)\rangle=\mathcal{C}$ as $h^{\prime}(x)$ is a nonzero polynomial. ∎
@@
$\tau(c(T_{\theta,\Im})(h^{\prime}))=c(x)h^{\prime}(x)=a(x)g(x)h^{\prime}(x)=a(x)h(x)g(x)=a(x)(x^{n}-1)=0$
for some $a(x)$ in $\mathbb{F}_{q}[x;\theta,\Im]/\langle x^{n}-1\rangle$ and $c(x)h^{\prime}(x)$ is taken modulo $x^{n}-1$. This implies $c(T_{\theta,\Im})(h^{\prime})=0$. Thus, $0=c(T_{\theta,\Im})(h^{\prime})=c_{0}+c_{1}T_{\theta,\Im}(h^{\prime})+c_{2}T_{\theta,\Im}^{2}(h^{\prime})+\cdots+c_{n-1}T_{\theta,\Im}^{n-1}(h^{\prime}).$ This shows that $(c_{0},c_{1},c_{2},\dots,c_{n-1}).H=0$ for any $c=(c_{0},c_{1},c_{2},\dots,c_{n-1})\in\mathcal{C}$. Also, $\tau(T_{\theta,\Im}^{k}(h^{\prime}))=x^{k}h^{\prime}(x)$ for $k=0,1,\dots,n-deg\leavevmode\nobreak\ h^{\prime}(x)-1=deg\leavevmode\nobreak\ g(x)-1$ and hence $\{h^{\prime},T_{\theta,\Im}(h^{\prime}),T_{\theta,\Im}^{2}(h^{\prime}),\dots,\\ T_{\theta,\Im}^{r-1}(h^{\prime})\}$ are linearly independent. ∎
@@
As $h^{\prime}(x)$ is a nonzero polynomial, we have $a(x)-c(x)b(x)g(x)=0$, which gives $a(x)=c(x)b(x)g(x)$. Therefore, $a(x)\in\mathcal{C}=\langle g(x)\rangle$. Thus, $\mathcal{C}^{\perp}\subseteq\mathcal{C}$. ∎
@@ 4 $(\gamma,\Delta)$-cyclic codes over $\mathscr{R}_{q,s}$
In this section, our main focus is to discuss the algebraic properties of $(\gamma,\Delta)$-cyclic codes over $\mathscr{R}_{q,s}$ via decomposition over $\mathbb{F}_{q}$. To do so, we consider a linear code $\mathcal{C}$ of length $n$ over $\mathscr{R}_{q,s}$. Towards this, we define
$\mathcal{C}_{i}=\left\{t_{i}\in\mathbb{F}_{q}^{n}\leavevmode\nobreak\ |\leavevmode\nobreak\ \sum_{i=0}^{s}\zeta_{i}t_{i}\in\mathcal{C},\leavevmode\nobreak\ \text{for some}\leavevmode\nobreak\ t_{0},t_{1},\dots,t_{i-1},t_{i+1},\dots,t_{s}\in\mathbb{F}_{q}^{n}\right\}$
for $0\leq i\leq s$. Then $\mathcal{C}_{i}$ is a linear code of length $n$ over $\mathbb{F}_{q}$ and $\mathcal{C}$ can be decomposed as
@@
Hence, $\Delta$ is a $\gamma$-derivation of $\mathscr{R}_{q,s}$. ∎
@@
Therefore, $\mathcal{C}$ is a $(\gamma,\Delta)$-cyclic code of length $n$ over $\mathscr{R}_{q,s}$. ∎
Theorem 4.3.
Let $\mathcal{C}=\displaystyle\bigoplus_{i=0}^{s}\zeta_{i}\mathcal{C}_{i}$ be a $(\gamma,\Delta)$-cyclic code of length $n$ over $\mathscr{R}_{q,s}$. Then $\mathcal{C}=\langle\zeta_{0}g_{0}(x),\zeta_{1}g_{1}(x),\dots,\zeta_{s}g_{s}(x)\rangle$ and $|\mathcal{C}|=q^{(s+1)n-\sum_{i=0}^{s}\deg(g_{i}(x))}$, where $g_{i}(x)$ is a generator polynomial of $\mathcal{C}_{i}$ for $i=0,1,2,\dots,s$.
Proof.
@@
On the other hand, we consider $\zeta_{0}f_{0}(x)g_{0}(x)+\zeta_{1}f_{1}(x)g_{1}(x)+\cdots+\zeta_{s}f_{s}(x)g_{s}(x)\in\langle\zeta_{0}g_{0}(x),\\ \zeta_{1}g_{1}(x),\dots,\zeta_{s}g_{s}(x)\rangle\subseteq\mathscr{R}_{q,s}[x;\gamma,\Delta]/\langle x^{n}-1\rangle$ where $f_{i}(x)\in\mathscr{R}_{q,s}[x;\gamma,\Delta]/\langle x^{n}-1\rangle$ for $i=0,1,2,\dots,s$. Then there exists $s_{i}(x)\in\mathbb{F}_{q}[x;\theta,\Im]/\langle x^{n}-1\rangle$ such that $\zeta_{i}f_{i}(x)=\zeta_{i}s_{i}(x)$ for $i=0,1,2,\dots,s$. This implies that $\langle\zeta_{0}g_{0}(x),\zeta_{1}g_{1}(x),\dots,\zeta_{s}g_{s}(x)\rangle\subseteq\mathcal{C}$. Thus, $\mathcal{C}=\langle\zeta_{0}g_{0}(x),\zeta_{1}g_{1}(x),\\ \dots,\zeta_{s}g_{s}(x)\rangle$. Moreover, $|\mathcal{C}|=|\mathcal{C}_{0}||\mathcal{C}_{1}|\cdots|\mathcal{C}_{s}|=q^{n-\deg(g_{0}(x))}q^{n-\deg(g_{1}(x))}\cdots q^{n-\deg(g_{s}(x))}\\ =q^{(s+1)n-\sum_{i=0}^{s}\deg(g_{i}(x))}$. ∎
Theorem 4.4.
@@
Conversely, let $\mathcal{C}^{\perp}\subseteq\mathcal{C}$ , then $\bigoplus_{i=0}^{s}\zeta_{i}\mathcal{C}_{i}^{\perp}\subseteq\displaystyle\bigoplus_{i=0}^{s}\zeta_{i}\mathcal{C}_{i}$. Now, considering modulo $\zeta_{i}$, we get $\mathcal{C}_{i}^{\perp}\subseteq\mathcal{C}_{i}$ for $i=0,1,2,\dots,s$. Thus, $h_{i}^{\prime}(x)h_{i}^{\prime}(x)$ is divisible by $x^{n}-1$ on the right for $i=0,1,2,\dots,s$. ∎